OpenStudy (anonymous):

how do you solve this equation 3=√b-1

OpenStudy (anonymous):

Add 1 to both sides to get\[4=\sqrt{b}\]and then square both sides to find b:\[(4)^2=(\sqrt{b})^2 \rightarrow 16=b\]